# Resource-Efficient Adaptation of Large Language Models for Text Embeddings via Prompt Engineering and Contrastive Fine-tuning

### v1

@@ Abstract
Large Language Models (LLMs) have become a cornerstone in Natural Language Processing (NLP), achieving impressive performance in text generation. Their token-level representations capture rich, human-aligned semantics. However, pooling these vectors into a text embedding discards crucial information. Nevertheless, many non-generative downstream tasks, like e.g. clustering, classification, or retrieval, still depend on accurate and controllable sentence- or document-level embeddings. We explore several adaptation strategies for pre-trained, decoder-only LLMs: (i) various aggregation techniques for token embeddings, (ii) task-specific prompt engineering, and (iii) text-level augmentation via contrastive fine-tuning. Combining these components yields state-of-the-art performance on the English clustering track of the Massive Text Embedding Benchmark (MTEB). An analysis of the attention map further shows that fine-tuning shifts focus from prompt tokens to semantically relevant words, indicating more effective compression of meaning into the final hidden state. Our experiments demonstrate that LLMs can be effectively adapted as text embedding models through a combination of prompt engineering and resource-efficient contrastive fine-tuning on synthetically generated positive pairs.
@@ 1 Introduction
Building on BERT, Stankevicius et al. summarize and evaluate numerous existing methods for Semantic Textual Similarity (STS), text clustering, and classification [8]. Their study explores information extraction from various layers, experiments with different pooling strategies, and applies post-processing techniques. Another notable method is Diagonal Attention Pooling (DITTO)[9], which addresses the bias of BERT-based sentence embeddings toward uninformative words. Their approach assigns weights to words based on model-derived importance scores and computes a weighted average of word representations from pretrained models, resulting in resource-efficient sentence embeddings.
@@
  • State-of-the-art performance: The resulting models achieve state-of-the-art accuracy on English clustering tasks in the MTEB, and the method generalizes well to classification tasks.
@@ 2 Related Work
As already mentioned, LLMs have recently become increasingly popular for text embedding due to their large-scale pretraining. BehnamGhader et al. proposed LLM2Vec, an unsupervised pipeline that creates an embedding model from any decoder-only LLM by enabling bidirectional attention, followed by masked token prediction and unsupervised contrastive learning [12]. By incorporating supervised learning, they achieved state-of-the-art performance on MTEB for a period. First, the authors enhanced the attention mechanism of the LLM by enabling bidirectional attention. They then retrained the model using masked token prediction to adapt it to this new attention scheme. Finally, the authors applied contrastive learning to embeddings obtained through mean-pooling of all attention outputs, fine-tuning the model directly in the embedding space to produce meaningful text representations. While mean-pooling yields the best results, the authors also experimented with using an end-of-sequence (EOS) token or weighted averaging.
Besides fine-tuning, prompting has also become increasingly popular. Jiang et al. use prompts to reshape the embedding space, focusing particularly on encouraging the model to condense a sentence into a single word — example given, "This sentence: "[X]" means in one word:“ [13]. They also experimented with in-context learning and subsequently applied contrastive learning. Building on this work, Zhang et al. also leverage LLMs and prompt engineering, testing new prompts and showing that, in a zero-shot inference setting, enforcing an explicit one-word limitation (EOL) improves performance, and the prompt choice significantly affects results [1]. When using efficient fine-tuning via contrastive learning on a human-annotated NLI dataset, they show that the prompt has little impact on STS tasks. They further introduce new zero-shot prompts, such as After thinking step by step, this sentence: "[X]" means in one word:“ [1], which they refer to Pretended Chain of Thought. They also present a second prompt called Knowledge Enhancement. Qwen3 Embedding has been recently released, comprising a family embedding and reranking models [14]. The models explicitly use an EOS token, leveraging its final hidden state as the text embedding. In their technical report, the authors describe the training pipeline as follows: (1) weakly supervised pretraining using large-scale synthetic paired data, (2) supervised fine-tuning with high-quality synthetic and labeled data, and (3) model merging using sampled checkpoints from the previous step [14]. The synthetic data was generated using the Qwen3-32B model, resulting in 150 million pairs for multi-task weakly supervised learning. Filtering the synthetic data by cosine similarity, they reduced it to 12 million high-quality supervised training data pairs. Their methods were evaluated on MTEB [10], achieving state-of-the-art performance.
@@ 3 Methods
Our methodology follows a three-stage workflow. First, we conduct a zero-shot evaluation of token aggregation strategies over pretrained LLM representations, using the English clustering subset of MTEB. Second, we shape the embedding space through prompting by wrapping clustering-oriented prompts around each input, encouraging representations that better capture text-level semantics. Third, we apply LoRA-based contrastive fine-tuning by adding low-rank adapters to the frozen backbone and optimizing them with a contrastive loss [15]. During this stage, we compare several positive-pair generation strategies to better align the embeddings produced under the selected prompt.
@@ 3.1 Embedding Extraction & Pooling
In zero-shot evaluation, EOS-token pooling consistently underperformed compared to the implicit last-token variant. As a result, all subsequent analyses focus on mean and last-token pooling. Mean pooling aggregates information across all positions, while last-token pooling leverages the auto-regressive objective, which encourages the final hidden state to encode the entire context. Although prior work reports competitive performance using explicit EOS pooling [1], our experiments indicate that this benefit does not generalize to the models and clustering tasks experimented in this study.
@@ 3.2 Prompt-Based Shaping of the Embedding Space
Prompt engineering is a proven technique for steering LLM representations in semantic textual similarity (STS) tasks [13, 1]. To adapt this idea to clustering, we re-evaluated the STS prompts and introduced cluster-specific variants. Table 1 lists the six prompts considered.
@@
We first replicate the three STS prompts — Explicit One-word Limitation (EOL), Pretended Chain of Thought (PCoT), and Summarize (SUM) — as introduced by [1]. We then design two clustering-oriented prompts, Cluster Command Wrapped (CCW), Cluster Command Prepended (CCP), and a question-style prompt (Question). CCP follows the instruction format used in Qwen3-Embedding [14], which takes the form {Instruction} {Query} <|endoftext|>.
@@ 3.3 Contrastive Fine-Tuning with Augmented Data
The third stage applies parameter-efficient LoRA fine-tuning to (i) adapt the model to the selected prompt and (ii) optimize it to produce the sentence embedding directly, rather than merely generating an intermediate representation for next-token prediction. Following [1], we attach LoRA adapters and train them using a contrastive loss in a self-supervised setting. As an initial source of positive pairs, we adopt the wiki1m_for_simcse dataset, a one-million-sentence Wikipedia corpus dataset used by SimCSE [5]. The motivation for using this dataset follows [12], which argues that since LLMs have already been trained on Wikipedia data, they will not acquire new domain knowledge, but only learn the fine-tuned technique. We then extended our study by applying a suite of natural-language data augmentation techniques [16, 17, 18], which are described in the remainder of the section.
@@
We progressively diversified the source of positive pairs as follows. First, we performed back-translation (English $\rightarrow$ German $\rightarrow$ English) on the Wikipedia corpus and fine-tuned a model on the resulting positive pairs. We then introduced lexical perturbations to increase surface-level variation. Next, we fine-tuned a model on a human-annotated NLI corpus previously used by Zhang et al. [1] and created by [19, 20, 21], which provides high-quality positive pairs. Finally, synthetic positive pairs were generated through LLM-based paraphrasing, using the prompts listed in Table 2, and used them in a separate fine-tuning stage.
@@ 4 Results
In this section, we address three questions: (i) Which pooling strategy produces the strongest zero-shot clustering performance? (ii) To what extent do clustering-oriented prompts improve these baselines? (iii) Which combination of fine-tuning, dataset, and augmentation strategy provides the best results? Unless stated otherwise, scores are averaged over the 21 English clustering datasets in MTEB v1.38.30. We also provide classification results to assess the generality of our findings. Following established MTEB protocol, we report single-run scores without confidence intervals or variance estimates to ensure direct comparability with prior work. Experiments were conducted on a single NVIDIA RTX A5000 (24 GB), except for models exceeding one billion parameters, which were trained using multiple A5000 GPUs. Our study focuses on two smaller LLMs, namely Llama-3.2-1B [22] and Qwen3-0.6B [23].
@@ 4.1 Embedding Extraction & Pooling
The zero-shot baseline was obtained by comparing three token aggregation strategies across Llama-3.2-1B and Qwen3-0.6B. As shown in Table 3, mean pooling delivers the highest average clustering accuracy, outperforming last-token pooling by approximately 12 % and EOS-token pooling by more than 20 % across both models. The Llama model also shows a consistent advantage over Qwen, reflecting its larger parameter count.
Model Mean-pooling Last-token EOS-token
Llama-3.2-1B 45.8 % 33.2 % 23.4 %
Qwen3-0.6B 39.8 % 28.9 % 22.8 %
@@ 4.2 Prompt-Based Shaping of the Embedding Space
Model Mean-pooling Last-token
@@
Table 4: Clustering accuracy with and without prompt-based input formatting. Results are reported using mean and last-token pooling across prompts. The CCW prompt leads to the largest gain when combined with last-token pooling, particularly for LLaMA-3.2-1B.
@@
On Llama-3.2-1B, the CCW prompt increases last-token accuracy from 33.2 % to 46.8 %, a gain of 13.6 percentage points without any additional inference cost. Other instruction-style prompts (EOL, PCoT) also improve performance, while the interrogative variant (Question) degrades it — suggesting that commands produce more consistent representations than questions. Qwen3-0.6B exhibits similar but smaller gains, as none of the tested prompts closes the gap to Llama — likely due to Qwen’s lower capacity. We were unable to identify an equally effective prompt for Qwen. The results further suggest that wrapping the prompt around the input may be beneficial. In a strictly zero-shot setting, we also evaluated larger checkpoints: Llama-3.2-3B, Qwen3-4B, and Qwen3-8B. Only the CCW prompt was used in this comparison, as shown in Table 5.
Model Mean-pooling Last-token

### v2

@@ Abstract
Large Language Models (LLMs) have become a cornerstone in Natural Language Processing (NLP), achieving impressive performance in text generation. Their token‑level representations capture rich, human‑aligned semantics. However, pooling these vectors into a text embedding discards crucial information. Nevertheless, many non‑generative downstream tasks, like e.g. clustering, classification, or retrieval, still depend on accurate and controllable sentence‑ or document‑level embeddings. We explore several adaptation strategies for pre‑trained, decoder‑only LLMs: (i) various aggregation techniques for token embeddings, (ii) task‑specific prompt engineering, and (iii) text‑level augmentation via contrastive fine‑tuning. Combining these components yields competitive performance on the English clustering track of the Massive Text Embedding Benchmark (MTEB). An analysis of the attention map further shows that fine‑tuning shifts focus from prompt tokens to semantically relevant words, indicating more effective compression of meaning into the final hidden state. Our experiments demonstrate that LLMs can be effectively adapted as text embedding models through a combination of prompt engineering and resource‑efficient contrastive fine‑tuning on synthetically generated positive pairs.
@@ 1 Introduction
Building on BERT, [8] summarize and evaluate numerous existing methods for Semantic Textual Similarity (STS), text clustering, and classification. Their study explores information extraction from various layers, experiments with different pooling strategies, and applies post-processing techniques. Another notable method is Diagonal Attention Pooling (DITTO)[9], which addresses the bias of BERT-based sentence embeddings toward uninformative words.
@@
  • Competitive performance: The resulting models achieve competitive accuracy on English clustering tasks on MTEB, and the method generalizes well to classification tasks.
@@
All source code, synthetic training data, and model checkpoints used in this study will be made publicly available.
@@ 2 Related Work
As already mentioned, LLMs have recently become increasingly popular for text embedding due to their large-scale pretraining. [12] proposed LLM2Vec, an unsupervised pipeline that creates an embedding model from any decoder-only LLM by enabling bidirectional attention, followed by masked token prediction and unsupervised contrastive learning. By incorporating supervised learning, they achieved state-of-the-art performance on MTEB for a period. First, the authors enhanced the attention mechanism of the LLM by enabling bidirectional attention. They then retrained the model using masked token prediction to adapt it to this new attention scheme. Finally, the authors applied contrastive learning to embeddings obtained through mean-pooling of all attention outputs, fine-tuning the model directly in the embedding space to produce meaningful text representations. Besides fine-tuning, prompting has also become increasingly popular. [13] use prompts to reshape the embedding space, focusing particularly on encouraging the model to condense a sentence into a single word — example given, "This sentence: "[X]" means in one word:“ [13]. Building on this work, [1] also leverage LLMs and prompt engineering, testing new prompts and showing that, in a zero-shot inference setting, enforcing an explicit one-word limitation (EOL) improves performance, and the prompt choice significantly affects results. When using efficient fine-tuning via contrastive learning on a human-annotated NLI dataset, they show that the prompt has little impact on STS tasks. They further introduce new zero-shot prompts, such as After thinking step by step, this sentence: "[X]" means in one word:“ [1], which they refer to Pretended Chain of Thought. They also present a second prompt called Knowledge Enhancement. Qwen3 Embedding has been recently released, comprising a family embedding and reranking models [14]. The models explicitly use an EOS token, leveraging its final hidden state as the text embedding. In their technical report, the authors describe the training pipeline as follows: (1) weakly supervised pretraining using large-scale synthetic paired data, (2) supervised fine-tuning with high-quality synthetic and labeled data, and (3) model merging using sampled checkpoints from the previous step [14]. The synthetic data was generated using the Qwen3-32B model, resulting in 150 million pairs for multi-task weakly supervised learning. Filtering the synthetic data by cosine similarity, they reduced it to 12 million high-quality supervised training data pairs. Their methods were evaluated on MTEB [10], achieving state-of-the-art performance.
@@ 3 Methods
Figure 1: Overview of the proposed pipeline. It comprises three stages: (A) data augmentation via back-translation and LLM paraphrasing, (B) LoRA fine-tuning with contrastive learning, and (C) benchmarking on MTEB with optional prompts, tuned adapters, and EOS tokens. Dashed boxes mark optional components.
Our methodology follows a three-stage workflow visualized in Figure 1. First, we conduct a zero-shot evaluation of token aggregation strategies over pretrained LLM representations, using the English clustering subset of MTEB. Second, we shape the embedding space through prompting by wrapping clustering-oriented prompts around each input, encouraging representations that better capture text-level semantics. Third, we apply LoRA-based contrastive fine-tuning by adding low-rank adapters to the frozen backbone and optimizing them with a contrastive loss [15]. During this stage, we compare several positive-pair generation strategies to better align the embeddings produced under the selected prompt.
@@ 3.1 Embedding Extraction & Pooling
In zero-shot evaluation, EOS-token pooling consistently underperformed compared to the implicit last-token variant. As a result, all subsequent analyses focus on mean and last-token pooling. Mean pooling aggregates information across all positions, while last-token pooling leverages the auto-regressive objective, which encourages the final hidden state to encode the entire context. Although prior work reports promising performance using explicit EOS pooling [1], our experiments indicate that this benefit does not generalize to the models and clustering tasks experimented in this study.
@@ 3.2 Prompt-Based Shaping of the Embedding Space
Prompt engineering is a proven technique for steering LLM representations in STS tasks [13, 1]. To adapt this idea to clustering, we re-evaluated the STS prompts and introduced cluster-specific variants. Table 1 lists the six prompts considered.
@@
We first replicate the three STS prompts — EOL, Pretended Chain of Thought (PCoT), and Summarize (SUM) — as introduced by [1]. We then design two clustering-oriented prompts, Cluster Command Wrapped (CCW), Cluster Command Prepended (CCP), and a question-style prompt (Question). CCP follows the instruction format used in Qwen3-Embedding [14], which takes the form {Instruction} {Query} <|endoftext|>.
@@ 3.3 Contrastive Fine-Tuning with Augmented Data
The third stage applies parameter-efficient LoRA fine-tuning to (i) adapt the model to the selected prompt and (ii) optimize it to produce the sentence embedding directly, rather than merely generating an intermediate representation for next-token prediction. Following [1], we attach LoRA adapters and train them using a contrastive loss in a self-supervised setting. As an initial source of positive pairs, we adopt the wiki1m_for_simcse dataset, a one-million-sentence Wikipedia corpus dataset used by SimCSE [5]. The motivation for using this dataset follows [12], which argues that since LLMs have already been trained on Wikipedia data, they will not acquire new domain knowledge, but only learn the fine-tuned technique. Negative pairs are generated by treating all non-positive samples in each training batch as negatives [16]. We then extended our study by applying a suite of natural-language data augmentation techniques [17, 18, 19], which are described in the remainder of the section.
@@
We progressively diversified the source of positive pairs as follows. First, we performed back-translation (English $\rightarrow$ German $\rightarrow$ English) on the Wikipedia corpus and fine-tuned a model on the resulting positive pairs. We then introduced lexical perturbations to increase surface-level variation. Next, we fine-tuned a model on a human-annotated NLI corpus previously used by [1] and created by [20, 21, 22], which provides high-quality positive pairs. Finally, synthetic positive pairs were generated through LLM-based paraphrasing, using the prompts listed in Table 2, and used them in a separate fine-tuning stage.
@@ 4 Results
In this section, we address three questions: (i) Which pooling strategy produces the strongest zero-shot clustering performance? (ii) To what extent do clustering-oriented prompts improve these baselines? (iii) Which combination of fine-tuning, dataset, and augmentation strategy provides the best results? Unless stated otherwise, scores are averaged over the 21 English clustering datasets in MTEB v1.38.30. We also provide classification results to assess the generality of our findings. In line with the established MTEB protocol, we report single-run scores without confidence intervals or variance estimates to ensure direct comparability with prior work. Because MTEB fixes dataset splits, random seeds, and evaluation scripts, repeated evaluations of the same model yield identical results. Experiments were conducted on a single NVIDIA RTX A5000 (24 GB), except for models exceeding one billion parameters, which were tested using multiple GPUs. Our study focuses on two smaller LLMs, namely Llama-3.2-1B [23] and Qwen3-0.6B [24].
@@ 4.1 Embedding Extraction & Pooling
The zero-shot baseline was obtained by comparing three token aggregation strategies across Llama-3.2-1B and Qwen3-0.6B. As shown in Table 3, mean pooling delivers the highest average clustering accuracy, outperforming last-token pooling by approximately 12 % and EOS-token pooling by around 20 % across both models. The Llama model also shows a consistent advantage over Qwen, reflecting its larger parameter count.
Model Mean‑pooling Last‑token EOS‑token
Llama‑3.2‑1B 45.8 % 33.2 % 23.4 %
Qwen3‑0.6B 39.8 % 28.9 % 22.8 %
@@ 4.2 Prompt-Based Shaping of the Embedding Space
Model Mean Last
@@
Table 4: Clustering accuracy with and without prompt-based input formatting. Results are reported using mean and last-token pooling across prompts. The CCW prompt leads to the largest gain when combined with last-token pooling.
@@
On Llama-3.2-1B, the CCW prompt increases last-token accuracy from 33.2 % to 46.8 %, a gain of 13.6 percentage points with barely any additional inference cost. Other instruction-style prompts (EOL, PCoT) also improve performance, while the interrogative variant (Question) degrades it — suggesting that commands produce more consistent representations than questions. Qwen3-0.6B exhibits similar but smaller gains, as none of the tested prompts closes the gap to Llama — likely due to Qwen’s lower capacity. We were unable to identify an equally effective prompt for Qwen. The results further suggest that wrapping the prompt around the input may be beneficial.
4.3 Contrastive Fine-Tuning with Augmented Data
Method Mean‑pooling
Llama‑3.2‑1B
Last‑token
Llama‑3.2‑1B
Mean‑pooling
Qwen3‑0.6B
Last‑token
Qwen3‑0.6B
Base 45.8 % 33.2 % 39.8 % 28.9 %
EOL 46.1 % 43.2 % 38.3 % 26.9 %
CCW 45.3 % 46.8 % 38.6 % 30.7 %
CCP 43.5 % 32.8 % 38.9 % 29.8 %
Base + nli 49.2 % 47.8 % 46.8 % 47.4 %
CCW + nli 52.5 % 52.2 % 49.2 % 49.6 %
CCP + nli 49.5 % 49.7 % 46.2 % 48.9 %
CCW + wiki‑translate 51.2 % 52.2 % 47.1 % 49.4 %
CCP + wiki‑translate 47.0 % 45.3 % 44.4 % 43.8 %
CCW + wiki‑translate + aug 51.7 % 52.6 % 46.8 % 47.7 %
CCP + wiki‑translate + aug 47.2 % 46.5 % 43.4 % 45.6 %
CCW + wiki‑prompt 52.7 % 54.5 % 48.8 % 50.9 %
CCP + wiki‑prompt 48.4 % 51.8 % 45.3 % 49.0 %
Consistent with prior findings, prompts benefit the last-token variant substantially more than mean pooling.
Training LoRA adapters on the human-curated NLI corpus without a prompt improves performance for both models, yet it does not surpass prompt-based fine-tuning. On Llama, the gap is particularly pronounced — fine-tuning NLI without a prompt achieves 47.8 %, barely outperforming CCW prompting alone at 46.8 %.
The best results are obtained using LLM-generated positive pairs, referred to as “wiki-prompt”. When combined with CCW, last-token accuracy reaches 54.5 % for Llama and 50.9 % for Qwen, representing the overall best scores. Mean pooling also performs strongly, peaking at 52.7 % for Llama on the same synthetic dataset. For Qwen, mean pooling achieves its best result on NLI data.
Detailed results for each task, including v_measure scores and standard deviations, are reported in Appendix B.2 for the CCW + wiki-prompt setting using Llama with last-token pooling. Also, all hyperparameter details are provided in Appendix A.
4.4 Comparison to other models on MTEB(eng, v2) Clustering
Table 6 compares our best model against a wide set of available embedding models. Proprietary systems such as gemini-embedding-001 [25] (59.39 %) and large-scale open models like gte-Qwen2-7B-instruct [26] (58.97 %) and Qwen3-Embedding-8B (58.57 %) currently define the state of the art. Slightly smaller variants such as Qwen3-Embedding-4B (57.51 %) and GritLM-7B [27] (50.82 %) further illustrate the performance advantages of scaling dedicated embedding models.
Within this landscape, our 1.3B-parameter checkpoint achieves 50.97 %, outperforming general-purpose models of comparable or smaller size such as GIST-large-Embedding-v0 [28] (48.84 %), multilingual-e5-large-instruct [29] (49.89 %), and NV-Embed-v2 [30] (47.66 %). Compared to the Qwen3-Embedding family, our approach trails the 4B and 8B variants, but remains close to Qwen3-Embedding-0.6B, despite being trained with far less compute and without task-specific prompting.
Importantly, this performance is achieved with a lightweight LoRA adaptation on Llama-3.2-1B, requiring under one GPU-hour on a single RTX A5000. This demonstrates that competitive clustering quality can be obtained without scaling to multi-billion-parameter embedding-specialized models, reinforcing the efficiency and adaptability of our method.
Model Param. Clustering
NV-Embed-v2 7.8B 47.66 %
GIST-large-Embedding-v0 0.3B 48.84 %
multilingual-e5-large-in. 0.6B 49.89 %
GritLM-7B 7.2B 50.82 %
gte-Qwen2-1.5B-instruct 1.5B 53.54 %
gte-Qwen2-7B-instruct 7.6B 58.97 %
Qwen3-Embedding-0.6B 0.6B 54.05 %
Qwen3-Embedding-4B 4B 57.51 %
Qwen3-Embedding-8B 8B 58.57 %
gemini-embedding-001 59.39 %
Our best (Llama-3.2-1B + CCW) 1.3B 50.97 %
Table 6: Comparison of clustering performance between our best Llama model, the Qwen3‑Embedding family, and other competitors. Results are reported for MTEB(eng, v2) Clustering, introduced by [31].
Model Mean Last
Llama-3.2-1B 72.5 % 68.1 %
Llama-3.2-1B + Prompt 73.0 % 70.3 %
Llama-3.2-1B + Prompt + FT 74.2 % 75.4 %
Qwen3-0.6B 70.2 % 63.0 %
Qwen3-0.6B + Prompt 70.3 % 64.1 %
Qwen3-0.6B + Prompt + FT 72.4 % 72.5 %
Table 7: Classification accuracy on the English subset of MTEB using the same configuration as in the clustering experiment. No hyperparameter adjustment was done. A classification-style prompt structurally aligned with the best-performing clustering prompt (CCW) was used.
4.5 Classification as a Downstream Task
The observed performance pattern closely mirrors that of the clustering tasks. In the zero-shot setting, mean pooling consistently outperforms last-token pooling, and Llama-3.2-1B maintains a slight advantage over Qwen3-0.6B. Introducing the CCW-style prompt yields modest improvements, again more pronounced for last-token embeddings. Fine-tuning with LoRA using the same hyperparameters as in the clustering experiments increases last-token accuracy to 75.4 % for Llama and 72.5 % for Qwen, establishing this as the best overall configuration. Notably, these improvements are obtained without classification-specific prompt tuning or hyperparameter optimization, highlighting the transferability of the pipeline across task types.
We further evaluate a key advantage of our method — its ability to incorporate domain-specific knowledge through targeted fine-tuning. To this end, we fine-tune Llama-3.2-1B on two MTEB classification datasets with labeled training splits: EmotionClassification and ToxicConversationsClassification[10]. Based on prior results showing its superiority in fine-tuned settings, we use last-token pooling exclusively. For each task, we construct a domain-specific contrastive training set by prompting an LLM to generate paraphrases for each sentence in the task’s training set, thereby creating positive pairs. Fine-tuning on these smaller classification datasets is even faster than in the clustering setting. As before, less than a full epoch is required, with performance typically peaking after approximately 100 optimization steps.
Method Emotion Toxic
Last 29.26 % 63.53 %
Last + Prompt 42.81 % 63.19 %
Fine-tuned (general) 45.33 % 67.38 %
Fine-tuned (task) 49.78 % 68.93 %
Table 8: Classification accuracy on EmotionClassification and ToxicConversationsClassification. “Last” denotes the last-token embedding without a prompt; “+ prompt” adds a classification-style prompt. “Fine-tuned (general)” refers to LoRA on the wiki-prompt corpus; “Fine-tuned (task)” uses LLM-generated positive pairs from each task’s training set.
Figure 2: Average attention received by each input token from the final token in the last transformer layer of Llama-3.2-1B, prompted with "This sentence: ‘a man is driving a car’ belongs to the following cluster:". Bars indicate values for at three stages of training: prompt-only, after 100 fine-tuning steps, and after 20,000 steps. Fine-tuning initially shifts attention to semantically meaningful tokens (“man”, “driving”, “car”), while prolonged training leads to a more uniform distribution.
Table 8 illustrates a consistent pattern. The classification prompt leads to a substantial improvement on the Emotion task but slightly deteriorates for Toxicity. Fine-tuning on the general wiki-prompt corpus enhances performance across both tasks. However, fine-tuning on the task-specific data achieves the highest scores, improving Emotion by 4.5 % and Toxicity by 1.6 % over the general model. While the extent of improvement varies, the addition of task-relevant supervision consistently results in further gains.
4.6 Analysis of Attention
To investigate how LoRA fine-tuning alters the model’s internal focus, we follow a method similar to that of [1]. Attention maps were generated for the probe sentence "a man is driving a car" and validated using additional sentences sampled from MTEB. Additional examples and detailed discussion are provided in Appendix B.3. The analysis examines the outgoing attention distribution of the final token in Llama-3.2-1B with the CCW prompt, comparing three settings illustrated in Figure 2: panel (a) shows the zero-shot model, panel (b) presents the best checkpoint after approximately 100 optimization steps, and panel (c) displays an over-trained checkpoint after 20,000 steps.
In all cases, the final token directs most of its attention to itself, a well-known behavior in causal transformers that preserves the accumulated context through residual connections. The key difference lies in the distribution of the remaining attention. After moderate fine-tuning, attention shifts away from the prompt tokens toward the content words man, driving, and car, emphasizing the sentence’s subject, predicate, and object — the elements most informative for downstream semantic tasks. With extended training, the focused attention dissipates; attention becomes more uniformly distributed, and model performance declines accordingly. These observations suggest that an appropriately timed fine-tuning phase encourages the model to encode task-relevant semantics into the final hidden state, while over-training diffuses this information and diminishes embedding quality.
5 Conclusion
This work focuses on turning pretrained LLMs into strong sentence-embedding models. We show that parameter-efficient fine-tuning, such as LoRA steered by a well-chosen prompt, can achieve high-quality embeddings. Additionally, sentence-level data augmentation like synthetic positive pairs enables promising contrastive self-supervised fine-tuning without manually labeled data. We raised Llama-3.2-1B and Qwen3-0.6B to highly competitive performance on the English clustering subset of MTEB, rivaling a dedicated embedding baseline in under an hour on a single RTX A5000. Attention maps indicate a more interpretable adaptation to downstream tasks.
Limitations
The current framework is limited to task-specific prompt formats like clustering or classification, English-only text data, and full-precision models. It does not yet incorporate task-agnostic prompt design, multilingual or cross-modal applications, or quantization methods for reducing resource consumption even further. Furthermore, the attention analysis presented here should not be interpreted as providing causal explanations; instead, it offers descriptive evidence. While the analysis cannot establish definitive cause-and-effect relationships, the consistent patterns observed across a broad range of test sentences increase confidence that the model tends to focus on the input to be encoded, with particular emphasis on the most relevant tokens.
Acknowledgements
We used language and writing support tools (ChatGPT, DeepL Write, and Grammarly) to improve readability and grammar. ChatGPT was also employed to assist with code generation and debugging. The authors remain solely responsible for the content of this work.
Appendix A Hyperparameter-Tuning
Table 9 lists the hyperparameters explored during fine-tuning. The final settings used in our experiments are highlighted in bold. We initially trained LoRA adapters on two attention projections, namely the query and value projections. However, we observed higher scores after extending optimization to include six projections in total. These additional projections also influence the feed-forward pathway and include the output, gate, up, and down projections. This broader parameter coverage is particularly well-suited for embedding tasks, as it allows the model to adapt both the attention mechanism and token-mixing components.
Hyperparameter Tested values
Dropout 0.05, 0.1, 0.2, 0.5
Learning rate $1\times 10^{-5}$, $\mathbf{5\times 10^{-5}}$, $5\times 10^{-4}$
Batch size 20, 60, 120, 500
Temperature 1, 0.5, 0.2, 0.1, 0.05
Projection layers 2, 6
All fine-tuning experiments were conducted on a single NVIDIA RTX A5000 GPU with 24 GB of memory. Peak validation performance typically emerged within 0.1 to 0.5 epochs, corresponding to approximately 100 optimization steps or 12,000 examples at the selected batch size of 120. In some cases, performance plateaued later, between 500 and 2,500 steps. To address this, we saved intermediate checkpoints and selected the best-performing model post hoc. In practice, fine-tuning runs completed under 30 minutes, demonstrating the method’s efficiency and suitability for low-resource environments. All additional hyperparameters can be found in the code in their final ideal state.
Appendix B Results
This appendix section provides supplementary experimental results and analyses that complement the findings in the main text. We report additional zero-shot comparisons across model sizes, detailed per-dataset performance with standard deviations, and further qualitative analyses such as attention visualizations.
B.1 Zero-Shot Comparison by Model Size
In a strictly zero-shot setting, we also evaluated larger checkpoints: Llama-3.2-3B, Qwen3-4B, and Qwen3-8B. Only the CCW prompt was used in this comparison, as shown in Table 10.
Model Mean Last